You are trying to decide whether to install a satellite dish or purchase a DVD player and rent DVD's. The satellite is a one time installation of $45. plus $34./month. To purchase the DVD player it is $185 and rent 8 DVD's per month at $2.50 each. At six months the satellite is a better deal and at 12 months the DVD is a better deal. What is the "Break even point"? or At how many months and at what cost will the two options be equal? Explain your answer.
Answer by jorel1380(3719)

45+34x=185+8(2.50)x
45+34x=185+20x
14x=140
x=10
At 10 months, the costs between the two options are the same. That's the "break even point". That's the point when the two options cost equal.